# Empirical Analysis of the Model-Free Valuation Approach: Hedging Gaps, Conservatism, and Trading Opportunities

### v1

@@ Abstract.
In this paper we study the quality of model-free valuation approaches for financial derivatives by systematically evaluating the difference between model-free super-hedging strategies and the realized payoff of financial derivatives using historical option prices from several constituents of the S&\&$500$ between $2018$ and $2022$.
@@ 1. Introduction
Building on the foundational works beiglbock2013model; beiglbock2016problem; henry2016explicit, which establish the connection between solutions of the martingale optimal transport problem, model-free price bounds for financial derivatives and associated hedging strategies, the literature on model-free approaches and their application to derivative valuation has grown rapidly and still continues to grow. See, for instance, ansari2024improved; cheridito2017duality; chiu2023model; davis2016model; dolinsky2014martingale; de2020bounds; fahim2016model; hobson2012robust; hobson2015robust; eckstein2021robust; lux2017improved, among many others. Using only available market data of liquidly traded put- and call options, the aforementioned model-free approaches compute price bounds for less liquidly traded (so called exotic) options solely under the paradigm of no arbitrage and of consistency with observed liquid option prices but, importantly, without using any assumptions on the stochastic dynamics of the underlying assets. The resultant price bounds are then calculated as extreme prices among all calibrated arbitrage-free models, and therefore can be understood to be model-free and to respect the so called Knightian uncertainty (knight1921risk), i.e., the unquantifiable risk of having chosen a wrong model.
@@
To the best of our knowledge there exists however little to no research providing empirical evidence for the actual conservativeness of model-free approaches, i.e., research answering the question how much the corresponding prices bounds are too wide in real financial markets, and whether they indeed prevent applicability. According to the duality theory developed in beiglbock2013model; beiglbock2016problem; henry2016explicit, to study the quality of model-free upper price bounds, one can equivalently study prices of model-free super-replication strategies (or sub-replication strategies for lower price bounds). Namely, if the gap between the realized payoff of a price bound-attaining model-free super-replication strategy and the realized derivative payoff is (too) large, this indicates that the strategy’s price and hence its initial cash endowment was too large as well and could be reduced to obtain a less conservative strategy with payoff closer to the derivative’s actual payoff.
@@
Relying on the above outlined idea, the present paper aims at closing an existing research gap by studying systematically the difference between model-free hedging strategies (corresponding to model-free price bounds) and realized payoffs of exotic derivatives. For our empirical study, we use historical option prices written on seven of the largest assets by market capitalization of the S&\&$500$ between $2018$ and $2022$. This paper is the first to empirically quantify the realized hedging performance and conservatism of model-free super-hedging strategies using real market data. Our findings challenge the prevailing common belief that model-free methods are impractically conservative, at least if the involved assets and sufficiently many call- and put options are liquidly traded in the market, as it is the case for our study.
@@
Our analysis shows that model-free hedging strategies lead to a hedging performance lacking not much behind the ones of industry-standard models such as the Black-Scholes-, Heston- and Dupire’s local volatility model. As, in contrast to the aforementioned industry-standard models, model-free super hedging strategies do not rely on a specific probabilistic model of the underlying security, they always lead to larger payoffs than the derivative given any realization of the security. We use this robustness property to propose a trading strategy that makes use of both our statistical description of the payoff gap and the model-free construction of the hedging strategy. Trading for this strategy is triggered whenever observed market quotes for the derivative are close enough to its model-free price bounds where the degree of proximity determines the probability of a positive payoff which can be exactly described according to our statistical analysis. The strategy is robust in the sense that up to the initial investment no additional money can be lost due to its model-free construction, consequently leading to a low-risk trading strategy with an easy-to-understand risk profile.
@@
The remainder of this paper is as follows. In Section 2 we provide an overview on the used data, Section 3 describes in detail the employed mathematical approach to compute the super-hedging strategies based on linear programming. Section 4 presents the results of our analysis when being applied to data, and Section 5 discusses and analyzes a trading strategy that is based on our statistical findings.
@@ 3.1. Model-free valuation of financial derivatives
$+\sum_{i=0}^{n-1}H_{i}(S_{t_{1}},\dots,S_{t_{i}})(S_{t_{i+1}}-S_{t_{i}}e^{r_{(t_{i+1}-t_{i})}\cdot(t_{i+1}-t_{i})}),$
where $r_{\tau}\in\mathbb{R}$ denotes the interest rate used for discounting of a time period of length $\tau$, $d\in\mathbb{R}$ denotes the cash position, $(\theta_{i,j}^{\rm call,+},\theta_{i,j}^{\rm call,-})\in\mathbb{R}_{\geq 0}^{2}$ the holdings of the call options, $K_{i,j}^{\rm call}\in\mathbb{R}_{\geq 0}$ the strikes of the call options, $M_{i}\in\mathbb{N}$ the number [^2] Note that we do not need to distinguish between the number of strikes considered for the call options since we can simply use the larger of the two and include options multiple times if necessary to match the number of options considered. of the call options with maturity $t_{i}$, $H_{i}(S_{t_{1}},\dots,S_{t_{i}})\in\mathbb{R}$ (while $H_{0}$ is a constant) denote the amount of the investment in the underlying security at time $t_{i}$.
@@
    As our analysis involves infrequent rebalancing of the dynamic positions $(H_{i})_{i}$, the effects of transaction costs for $(H_{i})_{i}$ are expected to be rather small, and thus, even though mathematically feasible, we have decided to not include assumptions on transaction costs related to trading in the underlying asset $S$ in our trading strategy. How to mathematically model semi-static trading that includes transaction costs is $S$ is described, e.g., in (cheridito2017duality, Section 3.1).
@@
(3.2) $\overline{\operatorname{P}}(\Phi)=\inf_{d,(\theta_{i,j}),(H_{i})}\Bigg{\{}$ $d+\sum_{i=1}^{n}\sum_{j=1}^{M_{i}}(\theta_{i,j}^{\rm call,+}h_{i,j}^{\rm call,+}-\theta_{i,j}^{\rm call,-}h_{i,j}^{\rm call,-})\leavevmode\nobreak\$
$\text{ subject to}\leavevmode\nobreak\ u_{d,(\theta_{i,j}),(H_{i})}(S)\geq\Phi(S)\text{ for all }S\in\mathbb{R}_{\geq 0}^{n}\Bigg{\}}$
@@
Note that if a financial model is calibrated to observed bid-ask prices of call options, then arbitrage-free prices of an exotic derivative $\Phi$ can be computed by $\mathbb{E}_{\mathbb{Q}}\left[e^{-r_{t_{n}-t_{0}}(t_{n}-t_{0})}\Phi(S_{t_{1}},\dots,S_{t_{n}})\right]$ where $\mathbb{Q}$ is a martingale measure that fulfils $\mathbb{E}_{\mathbb{Q}}\left[e^{-r_{t_{i}-t_{i}}(t_{i}-t_{0})}(S_{t_{i}}-K_{i,j}^{\rm call})^{+}\right]\in\left[h_{i,j}^{\rm{call},-},\leavevmode\nobreak\ h_{i,j}^{\rm{call},+}\right]$ for all $i,j$. If we denote the set of all such probability measures by $\mathcal{Q}$, then (cheridito2017duality, Section 3.1.3) or (neufeld2021model, Section A.1) show that under mild assumptions
@@ 3.2. Computation of model-free super-hedges with linear programming
$\text{ subject to}\leavevmode\nobreak\ u_{d,(\theta_{i,j}),(H_{0},H_{1})}(S)\geq\Phi(S)\text{ for all }S\in\mathbb{R}_{\geq 0}^{2}\Bigg{\}}.$
@@
$\text{ subject to }\leavevmode\nobreak\ u_{d,(\theta_{i,j}),(H_{0},\rm{spline}\left((H_{1}^{\ell})_{\ell}\right)}(S)\geq\Phi(S)\text{ for all }S\in\mathbb{G}\Bigg{\}}.$
@@ 4.1.1. Sub-hedging
In addition to the case of super-hedging discussed above, we also study the gap between realized outcomes of model-free sub-hedging strategies and the realized payoff of a forward start option with payoff $(S_{t_{2}}-S_{t_{1}})^{+}$. The statistics (median, quantile ranges) of the gap values under sub-hedging are illustrated in Figure 2. We remark that according to (3.6) the gap values in this situation are by construction mostly negative, but, in contrast to intuition, there are some mildly positive gap values in Figure 2 which we explain through the absence of transaction costs and due to the discretization inherent in Algorithm 1 in which we accept the result as long as the gap values do not exceed the tolerance level $\rm TOL$}.
@@ 4.3. Comparison with model-based hedges
In Figure  6 and Table 1, we compare the distribution of the resultant hedging errors (defined as in (3.6)) of the model-based hedging strategies with the hedging errors of the model-free super and sub-hedging strategies among all derivatives whose payoffs are $(S_{t_{2}}-S_{t_{1}})^{+}$ where the difference between $t_{1}$ and $t_{2}$ is fixed to $28$ days, and where the hedges are evaluated on the data described in Section 2.1. We observe that the hedges under all considered models lead to reasonably tight hedging gaps with interquartile ranges of [0.02,0.05],[0.03,0.03][-0.02,0.05],[-0.03,0.03], and [0.02,0.05][-0.02,0.05], respectively. The main observation is though that the interquartile ranges of the gaps of model-free super-hedges and sub-hedges are given by $[0.01,0.06]$, and $[0.00,0.04]$, respectively and therefore they are of similar size as the model-based hedges, contradicting the intuition and common belief that model-free approaches lead to overly conservative hedging outcomes. As by pricing-hedging duality (see, e.g., eberlein2008duality, hou2018robust), prices of hedges correspond to arbitrage-free prices of the corresponding derivative, the prices produced by model-free pricing approaches turn out to be competitive and rather close to prices produced by industry standard models (and thus traded prices of OTC derivatives) - an observation that we will use in the next section to explicitly construct a profitable trading strategy.
@@ 5.1. Description of the strategy
(5.1) $\min\{\mathfrak{M}(\Phi)-\underline{\operatorname{P}}(\Phi),\leavevmode\nobreak\ \leavevmode\nobreak\ \overline{\operatorname{P}}(\Phi)-\mathfrak{M}(\Phi)\}\leq c\cdot S_{0},$
@@
for a threshold $c=0.02$. We thus invest $\overline{C}=\overline{\rm P}(\Phi)-\mathfrak{M}(\Phi)=1.5$ in the model-free super-hedge, which dominates the derivative payoff path-wise. Assume that on maturity we observe $S_{t_{1}}=810$ and $S_{t_{2}}=870$, so that:

### v2

@@ Abstract.
In this paper we study the quality of model-free valuation approaches for financial derivatives by systematically evaluating the difference between model-free super-hedging strategies and the realized payoff of financial derivatives using historical option prices from several constituents of the S&\&P $500$ between $2018$ and $2022$.
@@ 1. Introduction
Robust and model-free pricing approaches date back at least to the seminal work by Hobson (hobson1998robust) who showed that lookback and related path-dependent options admit tight, model-independent price bounds and explicit super-replicating strategies based solely on observed call prices and the martingale property, without assuming any specific asset dynamics. More than a decade later, cox2011robust2 and cox2011robust established model-free upper and lower bounds on the arbitrage-free price of different types of Barrier options purely from observed vanilla option prices by constructing explicit super- and sub-hedging strategies and showing these bounds are tight via Skorokhod embedding techniques, all without assuming a specific probabilistic model for the underlying asset’s dynamics.
Based on cox2011robust, in obloj2010robust the authors analyzed how robust hedging strategies for digital double barrier options perform compared to traditional model-specific hedges (such as delta or delta/vega hedging). They find that under model uncertainty and transaction costs, robust hedges often achieve similar or better risk-adjusted performance and lower downside risk than classical methods across a range of market scenarios, albeit evaluated in hypothetical model-based forward markets.
In the 2010s, extending the above mentioned pathwise and Skorokhod-based pricing approaches, the literature on model-free derivative valuation developed rapidly. Building on the foundational works beiglbock2013model; beiglbock2016problem; henry2016explicit, which establish the connection between solutions of the martingale optimal transport problem, model-free price bounds for financial derivatives and associated hedging strategies, the literature on model-free approaches and their application to derivative valuation has grown rapidly and still continues to grow. See, for instance, ansari2024improved; cheridito2017duality; chiu2023model; davis2016model; dolinsky2014martingale; de2020bounds; fahim2016model; hobson2012robust; hobson2015robust; eckstein2021robust; lux2017improved, among many others. Using only available market data of liquidly traded put- and call options, the aforementioned model-free approaches compute price bounds for less liquidly traded (so called exotic) options solely under the paradigm of no arbitrage and of consistency with observed liquid option prices but, importantly, without using any assumptions on the stochastic dynamics of the underlying assets. The resultant price bounds are then calculated as extreme prices among all calibrated arbitrage-free models, and therefore can be understood to be model-free and to respect the so called Knightian uncertainty (knight1921risk), i.e., the unquantifiable risk of having chosen a wrong model.
@@
However, despite the extensive theoretical development of model-free pricing and the comparative performance analysis of robust hedges, apart from obloj2010robust, it remains largely unexplored how conservative model-free price bounds and the corresponding hedging strategies are when evaluated against realized payoffs in real financial markets.
According to the duality theory developed in beiglbock2013model; beiglbock2016problem; henry2016explicit, to study the quality of model-free upper price bounds, one can equivalently study prices of model-free super-replication strategies (or sub-replication strategies for lower price bounds).
Hence, empirical analysis of realized super-hedging payoffs provides direct information on the effective tightness of model-free price bounds in practice. Namely, if the gap between the realized payoff of a price-bound-attaining model-free super-replication strategy and the realized derivative payoff is persistently large across many observed price paths, this indicates that the strategy’s initial price, and hence its initial cash endowment, is overly conservative. In such cases, the initial capital could be reduced to obtain a less conservative strategy whose realized payoff is closer to the derivative’s actual payoff, with high probability under the empirical distribution induced by the observed paths.
This phenomenon is intrinsic to model-free hedging: super-replication strategies are constructed to dominate the derivative payoff pathwise, i.e., for every admissible price path. Since many of these admissible paths are highly unrealistic and effectively never occur in practice, the resulting hedges can appear excessively expensive when evaluated under empirically observed distributions.
@@
Relying on the above outlined idea, the present paper aims at closing an existing research gap by studying systematically the difference between model-free hedging strategies (corresponding to model-free price bounds) and realized payoffs of exotic derivatives. For our empirical study, we use historical option prices written on seven of the largest assets by market capitalization of the S&\&P $500$ between $2018$ and $2022$. This paper is, to the best of our knowledge, one of the first systematic empirical studies quantifying the realized hedging performance and conservatism of model-free super-hedging strategies using real market data. Our findings challenge the prevailing common belief that model-free methods are impractically conservative, at least if the involved assets and sufficiently many call- and put options are liquidly traded in the market, as it is the case for our study.
@@
Our analysis shows that model-free hedging strategies lead to a hedging performance comparable to that of industry-standard models such as the Black–Scholes, Heston, and Dupire local volatility models. As, in contrast to the aforementioned industry-standard models, model-free super-hedging strategies do not rely on a specific probabilistic model of the underlying security, they always lead to larger payoffs than the derivative given any realization of the security. From a practitioner’s perspective, this analysis suggests that model-free bounds may be sufficiently tight to support economically meaningful trading decisions. We use this robustness property to propose a trading strategy that makes use of both our statistical description of the payoff gap and the model-free construction of the hedging strategy. Trading for this strategy is triggered whenever observed market quotes for the derivative are close enough to its model-free price bounds where the degree of proximity determines the probability of a positive payoff which can be exactly described according to our statistical analysis. The strategy is robust in the sense that up to the initial investment no additional money can be lost due to its model-free construction, consequently leading to a low-risk trading strategy with an easy-to-understand risk profile.
While obloj2010robust study the performance of robust hedging strategies relative to classical hedges in hypothetical model-based markets, our paper provides, to the best of our knowledge, the first systematic empirical analysis of the conservativeness of model-free price bounds themselves using real market data, and leverages this insight to construct a provably low-risk trading strategy.
@@
The remainder of this paper is as follows. In Section 2 we provide an overview of the data used, Section 3 describes in detail the employed mathematical approach to compute the super-hedging strategies based on linear programming. Section 4 presents the results of our analysis when being applied to data, and Section 5 discusses and analyzes a trading strategy that is based on our statistical findings.
@@ 3.1. Model-free valuation of financial derivatives
${+\sum_{i=0}^{n-1}e^{r_{(t_{n}-t_{0})}\cdot(t_{n}-t_{0})}\left(H_{i}(S_{t_{1}},\dots,S_{t_{i}})(\widetilde{S}_{t_{i+1}}-\widetilde{S}_{t_{i}})-\varepsilon|\Delta H_{i}|\right),}$
where $r_{\tau}\in\mathbb{R}$ denotes the interest rate used for discounting of a time period of length $\tau$, and where $\widetilde{S}_{t_{i}}:=S_{t_{i}}e^{-r_{(t_{i}-t_{0})}\cdot(t_{i}-t_{0})}$ denotes the discounted stock price. Moreover, $d\in\mathbb{R}$ denotes the cash position, $(\theta_{i,j}^{\rm call,+},\theta_{i,j}^{\rm call,-})\in\mathbb{R}_{\geq 0}^{2}$ the holdings of the call options, $K_{i,j}^{\rm call}\in\mathbb{R}_{\geq 0}$ the strikes of the call options, $M_{i}\in\mathbb{N}$ the number [^2] Note that we do not need to distinguish between the number of strikes considered for the call options since we can simply use the larger of the two and include options multiple times if necessary to match the number of options considered. of the call options with maturity $t_{i}$, $H_{i}(S_{t_{1}},\dots,S_{t_{i}})\in\mathbb{R}$ (while $H_{0}$ is a constant) denote the amount of the investment in the underlying security at time $t_{i}$. The expression $\varepsilon|\Delta H_{i}|$ with $|\Delta H_{i}|:=|H_{i}(S_{t_{1}},\dots,S_{t_{i}})-H_{i-1}(S_{t_{1}},\dots,S_{t_{i-1}})|$ is accounting for per-share transaction costs $\varepsilon>0$ incurred from changing the position $H_{i}$, where $H_{-1}\equiv 0$. In our empirical implementation, we use $\varepsilon=0.0035$ USD per share [^3] This choice of $\varepsilon$ is implied by interactivebrokers. According to interactivebrokers the per share transaction costs are not larger than $0.0035\mathdollar$..
@@
    Note that the trading dates $(t_{i})_{i=1,\dots,n}$ of the dynamic strategy in (3.1) coincide exactly with the expiration dates of the considered call options. This is a natural construction, as allowing for more trading dates beyond the expiration dates, does not change the resultant price bounds computed in (3.2), see also (henry2017model, Corollary 2.1) and (sester2023intermediate, Lemma 4.2).
@@
(3.2) $\overline{\operatorname{P}}(\Phi)=\inf_{d,(\theta_{i,j}),(H_{i})}\Bigg\{$ $d+\sum_{i=1}^{n}\sum_{j=1}^{M_{i}}(\theta_{i,j}^{\rm call,+}h_{i,j}^{\rm call,+}-\theta_{i,j}^{\rm call,-}h_{i,j}^{\rm call,-})\penalty 10000\$
$\text{ subject to}\penalty 10000\ u_{d,(\theta_{i,j}),(H_{i})}(S)\geq\Phi(S)\text{ for all }S\in\mathbb{R}_{\geq 0}^{n}\Bigg\}$
@@
Throughout the paper, the target payoff $\Phi(S_{t_{1}},\dots,S_{t_{n}})$ is specified in undiscounted currency units at maturity $t_{n}$. For this reason, the superhedging condition
$u_{d,(\theta_{i,j}),(H_{i})}(S_{t_{1}},\dots,S_{t_{n}})\;\geq\;\Phi(S_{t_{1}},\dots,S_{t_{n}})$
is formulated in undiscounted terms, so that both sides of the inequality are expressed in the same units. This choice is without loss of generality. Introducing the deterministic discount factors $B(t)=e^{r_{(t-t_{0})}(t-t_{0})}$ and the discounted asset $\tilde{S}_{t}:=S_{t}/B(t)$, the above inequality is equivalent to the discounted condition
$\tilde{u}(\tilde{S})\;\geq\;\tilde{\Phi}(\tilde{S}),\qquad\tilde{u}:=\frac{u}{B(t_{n})},\quad\tilde{\Phi}:=\frac{\Phi}{B(t_{n})}.$
Hence, the discounted and undiscounted formulations differ only by a deterministic change of numéraire. For clarity, we may equivalently write the discounted strategy payoff as
$\tilde{u}_{d,(\theta_{i,j}),(H_{i})}$ $=d+\sum_{i=1}^{n}\sum_{j=1}^{M_{i}}(\theta^{\mathrm{call},+}_{i,j}-\theta^{\mathrm{call},-}_{i,j})\cdot e^{-r_{(t_{i}-t_{0})}(t_{i}-t_{0})}(S_{t_{i}}-K^{\mathrm{call}}_{i,j})_{+}\,$
$\qquad+\sum_{i=0}^{n-1}\left(H_{i}(S_{t_{1}},\dots,S_{t_{i}})\big(\tilde{S}_{t_{i+1}}-\tilde{S}_{t_{i}}\big)-\varepsilon|\Delta H_{i}|\right).$
[…]
Remark 3.3.
Note that if a financial model is calibrated to observed bid-ask prices of call options, then arbitrage-free prices of an exotic derivative $\Phi$ can be computed by $\mathbb{E}_{\mathbb{Q}}\left[e^{-r_{t_{n}-t_{0}}(t_{n}-t_{0})}\Phi(S_{t_{1}},\dots,S_{t_{n}})\right]$ where $\mathbb{Q}$ is a martingale measure that fulfils $\mathbb{E}_{\mathbb{Q}}\left[e^{-r_{t_{i}-t_{i}}(t_{i}-t_{0})}(S_{t_{i}}-K_{i,j}^{\rm call})^{+}\right]\in\left[h_{i,j}^{\rm{call},-},\penalty 10000\ h_{i,j}^{\rm{call},+}\right]$ for all $i,j$. If we denote the set of all such probability measures by $\mathcal{Q}$, then (cheridito2017duality, Section 3.1.3) or (neufeld2021model, Section A.1) show that under mild assumptions
@@ 3.2. Computation of model-free super-hedges with linear programming
$\text{ subject to}\penalty 10000\ u_{d,(\theta_{i,j}),(H_{0},H_{1})}(S)\geq\Phi(S)\text{ for all }S\in\mathbb{R}_{\geq 0}^{2}\Bigg\}.$
@@
$\text{ subject to }\penalty 10000\ u_{d,(\theta_{i,j}),(H_{0},\rm{spline}\left((H_{1}^{\ell})_{\ell}\right)}(S)\geq\Phi(S)\text{ for all }S\in\mathbb{G}\Bigg\}.$
@@ 4.1.1. Sub-hedging
In addition to the case of super-hedging discussed above, we also study the gap between realized outcomes of model-free sub-hedging strategies and the realized payoff of a forward start option with payoff $(S_{t_{2}}-S_{t_{1}})^{+}$. The statistics (median, quantile ranges) of the gap values under sub-hedging are illustrated in Figure 2. We remark that according to (3.6) the gap values in this situation are by construction mostly negative, but, in contrast to intuition, there are some mildly positive gap values in Figure 2 which we explain through the discretization inherent in Algorithm 1 in which we accept the result as long as the gap values do not exceed the tolerance level $\rm TOL$}.
@@ 4.3. Comparison with model-based hedges
In Figure 6 and Table 1, we compare the distribution of the resultant hedging errors (defined as in (3.6)) of the model-based hedging strategies with the hedging errors of the model-free super and sub-hedging strategies among all derivatives whose payoffs are $(S_{t_{2}}-S_{t_{1}})^{+}$ where the difference between $t_{1}$ and $t_{2}$ is fixed to $28$ days, and where the hedges are evaluated on the data described in Section 2.1. We observe that the hedges under all considered models lead to reasonably tight hedging gaps with interquartile ranges of [0.02,0.05],[0.03,0.03][-0.02,0.05],[-0.03,0.03], and [0.02,0.05][-0.02,0.05], respectively. The main observation is though that the interquartile ranges of the gaps of model-free super-hedges and sub-hedges are given by $[0.01,0.06]$, and [0.04,0.00]-[0.04,0.00], respectively and therefore they are of similar size as the model-based hedges, contradicting the intuition and common belief that model-free approaches lead to overly conservative hedging outcomes. As by pricing-hedging duality (see, e.g., eberlein2008duality, hou2018robust), prices of hedges correspond to arbitrage-free prices of the corresponding derivative, the prices produced by model-free pricing approaches turn out to be competitive and rather close to prices produced by industry standard models (and thus traded prices of OTC derivatives) - an observation that we will use in the next section to explicitly construct a profitable trading strategy.
@@ 5.1. Description of the strategy
(5.1) $\min\{\mathfrak{M}(\Phi)-\underline{\operatorname{P}}(\Phi),\penalty 10000\ \penalty 10000\ \overline{\operatorname{P}}(\Phi)-\mathfrak{M}(\Phi)\}\leq c\cdot S_{0},$
@@
for a threshold $c=0.02$. We thus invest $\overline{C}=\overline{\rm P}(\Phi)-\mathfrak{M}(\Phi)=1.5$ by selling the option and buying the model-free super-hedge, which dominates the derivative payoff path-wise. Assume that on maturity we observe $S_{t_{1}}=810$ and $S_{t_{2}}=870$, so that: